OpenStudy (anonymous):

Find the volume V of the solid obtained by rotating the region bounded by the given curves about the specified line. y = ln 3x, y = 4, y = 6, x = 0; about the y-axis I cant seem to get the integration right, here is what I have so far: y=ln3x e^y=3x 1/3 e^y=x pi/3 integral from 4 to 6 of (e^y)^2 I checked my answer on wolfram, and got the same integral, so it must be a problem with my setup.

OpenStudy (anonymous):

Hey Shall I ?

OpenStudy (anonymous):

Please, I'm still stuck

OpenStudy (anonymous):

Well I don't know integration but see this This is from yahoo I would be glad if it helps

OpenStudy (anonymous):

Well if you think about it; the clue is in the axis of rotation. It's rotated around the y-axis and not the x-axis as usual; as well as this; the limits of integration are y-values; so this should tell you that a function of y needs of be integrated; not a function of x. So instead of π∫ y^2 dx; the formula we have to use is; π∫ x^2 dy So now we need x in terms of y. Easy enough: y = ln(3x) e^y = 3x x = (1/3)e^y So then x^2 = (1/9)e^2y So what we're integrating here is: π∫ x^2 dy = π∫ (1/9)e^2y dy between the limits of 2 and 6. I'm sure you can mange that :)

OpenStudy (anonymous):

Ah, I think I loved the 1/3 out of the integral before squaring it.

OpenStudy (anonymous):

Hey I have a site for u to practice do u want it ?

OpenStudy (anonymous):

Yeah, sure

OpenStudy (anonymous):

http://mathpost.asu.edu/~zinzer/266HW7_Sol.pdf

OpenStudy (anonymous):

Nice! thanks

OpenStudy (anonymous):

Welcome

OpenStudy (anonymous):

Hey thank u for givin me a medal do u understand it now ?

OpenStudy (anonymous):

Yeah, it was a careless mistake, I got it right now. I pulled out the 1/3 before i squared it.

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

And the site?

OpenStudy (anonymous):

It it useful.

OpenStudy (anonymous):

r u from india

OpenStudy (anonymous):

No, USA

OpenStudy (anonymous):

ok